14C decays by emitting a β- particle. This decay process has a half-life of 5.72×103 yr. Answer the following questions about this process and report all answers to three significant figures.  

A. Determine the initial activity (in dis/yr) of a 5.81 mg sample of this isotope.

B. Determine the time (in yr) that it takes for the activity of 14C to decrease to 8.39×107 dis/yr.
Answer #1

half life time is 5.72 \times 10^{3} yr

decay constant will be:-

\lambda = 0.693/5.72 \times 10^{3}

\lambda = 1.21 \times 10^{-4}

A) initial activity will be given as:-

A=\lambda \times N

A= activity

N = number of undecayed nuclei

here N = 5.81\times 6.023\times 10^{23}/14

=24.9\times 10^{19}

puttinG values of both we get:-

\lambda =24.9\times 10^{19}\times 1.21\times 10^{-4}

\lambda =30.1\times 10^{15} dis/yr

B) the time can be calculated as:-

8.39 \times 10^{7}= 30.1\times 10^{15}\times e^{(-1.21 \times 10^{-4})\times t}

2.78 \times 10^{-9}= e^{(-1.21 \times 10^{-4})\times t}

taking natural log both sides we have :-

ln(2.78) -9\times 2.303= {(-1.21 \times 10^{-4})\times t}

1.02 -20.7= {(-1.21 \times 10^{-4})\times t}

t= 1.63 \times 10^{5} Yr